OpenStudy (anonymous):

Rewrite with only sin x and cos x. cos 2x + sin x

OpenStudy (anonymous):

The double angle formula for cosine is cos squared x- sin squared x. There are other ways to write the formula if you need a different answer. Do you think you can work with that?

OpenStudy (calculusfunctions):

Do you know the identity for cos 2x. There are two more ways to express cos 2x other than what @L.T. has suggested but I don't just want to give you the answer. I want to teach you how to arrive a the answer on your own. Would you like my help?

OpenStudy (anonymous):

yes please

OpenStudy (calculusfunctions):

@needinghelpyo I was helping someone else and just noticed your reply. Are you on line right now?

OpenStudy (anonymous):

yes

OpenStudy (calculusfunctions):

OK to start with do you understand what myself and @L.T. said above and do you know where to take it from there?

OpenStudy (anonymous):

no i dont know what the double angle area is its an online class and i dont understand the lesson for this segement

OpenStudy (calculusfunctions):

You mean the double angle identities. If you don't know them then here they are. You should know them by heart. \[\cos 2\theta =\cos ^{2}\theta -\sin ^{2}\] \[\cos 2\theta =2\cos ^{2}\theta -1\] \[\cos 2\theta =1-2\sin ^{2}\theta\] Now how would you use these to answer your question. Any idea?

OpenStudy (calculusfunctions):

*that is \[\sin ^{2}\theta\]at the end of the first one.

OpenStudy (anonymous):

would the answer be 1+2 sin^2x +sinx?

OpenStudy (calculusfunctions):

No, if the answer is to be written purely in terms of sin x. Close but it seems like you made an error somewhere. If you can show me your work then I can help you understand where you went wrong.

OpenStudy (anonymous):

i dont know how but then the other one i got is 1+3sin^2x

OpenStudy (calculusfunctions):

Didn't you say that you have to write one answer in terms of sine and the other in terms of cosine?

OpenStudy (calculusfunctions):

Can you show me your work for the first answer as I asked please?